The amounts of nicotine in a certain brand of cigarette are normally distributed with a mean of 0.918 g and a standard deviation of 0.304 g. The company that produces these cigarettes claims that it has now reduced the amount of nicotine. In what range would you expect to find the middle 80% of amounts of nicotine in these cigarettes (assuming the mean has not changed)?

(Enter your answers in ascending order...smaller on left, larger on right. Also, enter your answers accurate to four decimal places.)

If you were to draw samples of size 50 from this population, in what range would you expect to find the middle 80% of most average amounts of nicotine in the cigarettes in the sample?


(Enter your answers in ascending order...smaller on left, larger on right. Also, enter your answers accurate to four decimal places.)
Answer #1

The population mean being known, the confidence interval is obtained using a Z-distribution (Gaussian curve) of the mean based on standard deviation.

The critical Z-score for a two-tail 80% confidence level, viz. 0.2 signifiacne level is

Z_{\alpha /2}=Z_{0.1}=1.28155

The confidence interval is obtained as

\mu -Z_{\alpha /2}\cdot \sigma <\overline{x}<\mu +Z_{\alpha /2}\cdot \sigma

0.918-1.28155\times 0.304 <\overline{x}<0.918+1.28155\times 0.304

0.5284 <\overline{x}<1.3076

For a sample size of 50, we need to consider standard error for mean instead of standard deviation

\textup{Standard Error, }\sigma _{\overline{x}}=\frac{\sigma }{\sqrt{n}}=\frac{0.304}{\sqrt{50}}=0.0424264

The confidence interval of sample mean is obtained as

\mu -Z_{\alpha /2}\cdot \sigma_{\overline{x}} <\overline{x}<\mu +Z_{\alpha /2}\cdot \sigma_{\overline{x}}

0.918-1.28155\times 0.0424264<\overline{x}<0.918+1.28155\times 0.0424264

0.8636<\overline{x}<0.9724